OpenStudy (nicole143):

How do you find the quartic function that has its only two real zeros as x = -1 and x= -3? @mathmale

OpenStudy (solomonzelman):

you can write the roots as (x+1) and (x+3) f(x)=(x+1)(x+3) expand the parenthesis, and you get f(x)=x^2+4x+3

OpenStudy (nicole143):

Do you write it with q(x) on the end as well? @SolomonZelman

OpenStudy (solomonzelman):

Doesn't really matter, it's q(x) or f(x) q(x) or f(x) is just saying it's a function

OpenStudy (nicole143):

So to solve it out it would read f(x) = (x+1)(x+3)(x+1)?

OpenStudy (nicole143):

Then you would factor and simplify?

OpenStudy (solomonzelman):

yeah, but it's f(x)=(x+1)(x+3) without the last (x+1)

OpenStudy (nicole143):

But that doesn't give me enough to work with. What you said doesn't match my choices.

OpenStudy (nicole143):

I found the answer but I just substituted the -1 and -3 into all of the equation. I still want to know how to do it if I were to not have multiple choice.

OpenStudy (nicole143):

@crystalrock ?

OpenStudy (anonymous):

yea?

OpenStudy (nicole143):

Do you know how to solve?

OpenStudy (anonymous):

no, sorry. I am horrible at math

OpenStudy (nicole143):

Oh, okay..

OpenStudy (nicole143):

@nincompoop Could you help?

jhonyy9 (jhonyy9):

do you know this formula x^2 +Sx +P =0 where S = x_1 +x_2 and P = x_1 *x_2 for x_1 and x_2 roots of an quadratic

jhonyy9 (jhonyy9):

do you understand it ?

OpenStudy (nicole143):

I don't believe I do.. I may have seen it a little differently though

jhonyy9 (jhonyy9):

so x_1 = -1 and x_2 = -3 continue it please

OpenStudy (nicole143):

x^2 + x(x _ 1 + x _ 2) + x_1 * x_2 ??

OpenStudy (nicole143):

Before we do this may I ask you a quick question about a different problem?

jhonyy9 (jhonyy9):

sorry i have missed an ---- there is minus Sx the midle term so x^2 -Sx +P =0

OpenStudy (nicole143):

May I ask my other question?

jhonyy9 (jhonyy9):

ok i will try answer it ,than i can

OpenStudy (nicole143):

Can you find the two real zeros for an equation when the equation doesn't have an "a"?

OpenStudy (nicole143):

All real zeros - not two.. sorry.

OpenStudy (nicole143):

Or do you just make "a" 1 because there is nothing there?

jhonyy9 (jhonyy9):

how you think it ? because a exist allways when not is there nothing so than a = 1 yes

OpenStudy (nicole143):

Okay so if you had and equation like this; y = (x-h) + k would would make it into this; y = 1(x-h) + , right?

OpenStudy (nicole143):

+k on the end of the second equation - sorry

jhonyy9 (jhonyy9):

yes but why is this importantly ?

OpenStudy (nicole143):

The question on general or involving the equation?

OpenStudy (nicole143):

I had another question I had to do but hand't know to simply put 1 as "a" when "a" was missing.

jhonyy9 (jhonyy9):

write it here now please

OpenStudy (nicole143):

Okay, this is the equation dealing with my question about "a".

OpenStudy (nicole143):

Is that okay?

jhonyy9 (jhonyy9):

yes i think is right but we dont make never nothing with this 1 not is necessary writing there wnen you wann solve an exercise like this

OpenStudy (nicole143):

Oh, okay. Can I leave it though?

jhonyy9 (jhonyy9):

because you know sure that 7/1 = 7

OpenStudy (nicole143):

Yes, okay. I wrote the last answer as x = 3square root 7 + 12

OpenStudy (nicole143):

Thank you for your help.

jhonyy9 (jhonyy9):

ok but attention this not is square root this has a name like cuberoot

jhonyy9 (jhonyy9):

because squarroot has indice 2

OpenStudy (nicole143):

Oh i did not know that, thank you.

jhonyy9 (jhonyy9):

yw my pleasure good luck bye

OpenStudy (nicole143):

Bye. And thanks again.

OpenStudy (mathmale):

How do you find the quartic function that has its only two real zeros as x = -1 and x= -3? I think you mean "quadratic function." If the function is truly quadratic, its highest x power is x^2. We want to find the quadratic function that has real zeros {-1,-3). Represent this function as y=(x-[-1])(x-[-3]). That reduces to y=(x+1)(x+3). Expand this using the "foil" method. Write the result in the form y=ax^2 + bx + c.

OpenStudy (nicole143):

Hello, @mathmale thank you for this. I got what I need for that problem but if you should think of a trick to easily find factors of 400 that equal -40 I would be very appreciative! I can't seem get the right numbers.

OpenStudy (mathmale):

Hi, Nicole! If these two factors of 400 equal -40 when added together, ... Nicole, did you want to finish this problem or work on a new one? it's up to you.

OpenStudy (nicole143):

I would like to work a new one. I'll open a new post.

OpenStudy (mathmale):

Great. Please do!